Let T: P2 → P2 be defined by T(p(x)) p(7x + 8) (a) Find the determinant of T. (b) Find the eigenvalues of T. :
Answer #1

a we Know that B= B={1,2,42} is basis of IP then T(1) = 1 [co] B o o 8 T() = 7x+8 Stear]. 7 T(R2) = (5x+ 8) = 64+ 112x + 49x